Two contractors were working together to finish concrete. One can finish 25 sqaure yards in 4 hours, and the other can finish 25 sqaure yards in 5 hours. How much concrete can they finish in 8 hours working together?
Answer by jorel1380(3719)

1/4+1/5=1/t
5/20+4/20=9/20=1/t
9t=20
t=20/9
The two contractors can finish 25 square yards in 20/9 hours
8/(20/9)=72/20=3.6
3.6x25=90 sq. yds..